OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 18x + 71

OpenStudy (anonymous):

it is going to look like \(y=(x+9)^2+k\) for some \(k\) to find \(k\) replace \(x\) by \(-9\) in the original expression and see what you get because \(y=(-9+9)^2+k=0+k=k\)